# Learning to Adapt through Bio-Inspired Gait Strategies for Versatile Quadruped Locomotion

### v2

@@ Abstract
Deep reinforcement learning (DRL) has revolutionised quadruped robot locomotion, but existing control frameworks struggle to generalise beyond their training-induced observational scope, resulting in limited adaptability and gait proficiency. In contrast, animals achieve exceptional adaptability through gait transition strategies, diverse gait utilisation, and seamless adjustment to immediate environmental demands. Inspired by these capabilities, we present a novel DRL framework that incorporates key attributes of animal locomotion: gait transition strategies, pseudo gait procedural memory, and adaptive motion adjustments. This approach enables our framework to achieve unparalleled adaptability, demonstrated through blind zero-shot deployment on complex terrains and recovery from critically unstable states. Our findings offer valuable insights into the biomechanics of animal locomotion, paving the way for robust, adaptable robotic systems.
@@
Originally inspired by the remarkable adaptability of quadruped mammal locomotion—an ability shaped by innate and environmentally induced factors [1, 2, 3]—the field of quadruped robotics has made significant efforts to develop equally proficient locomotion frameworks. Presently, the most advanced systems rely on end-to-end deep reinforcement learning (DRL), which involves training a multilayer perceptron (MLP) [4] capable of navigating diverse environments.
@@
These frameworks demonstrate impressive realisation of terrestrial locomotion skills, which can be classified into two groups (further detailed in Supplementary Information Section 1); Froude-characterised locomotion (no desired velocity normal to the ground plane and hence upholds the assumptions of the Froude number [5]) such as walking or running in real-world [6, 7], urban [8, 9] and deformable[7] terrains, and Froude-free locomotion (features a desired velocity normal to the ground plane) such as jumping between platforms [10], climbing over obstacles [11], and sure-footedness [12]. Despite these achievements, when it comes to Froude-characterised locomotion, which can account for $\approx 70-90\%$ daily animal locomotion [13, 14], their adaptability remains constrained, as most systems are limited to deploying a single targeted gait or locomotion strategy.
@@
Figure 1: Instilling the core animal locomotion proficiency attributes within a DRL locomotion framework. From taking an abstracted view of animal locomotion to determine the attributes of proficient locomotion, we instil animal gait transition strategies within the gait selection policy, $\pi_{G}$, gait procedural memory is embedded within the bio-inspired gait scheduler, and adaptive motion adjustments are realised by the locomotion policy, $\pi_{L}$. $\pi_{G}$ has been trained to minimse the animal gait transition metrics applied to the quadruped robot based on the current robot state, $\bm{s}$, and relevant bio-inspired gait scheduler output, $\bm{\beta}_{G}$, to select the optimal gait, $\Gamma^{*}$. The bio-inspired gait scheduler then generates the gait references informed by $\bm{s}$ from encoded high-level gait parameters. The gait references, $\bm{\beta}_{L}$, are then passed to $\pi_{L}$ to inform of any adjustments to the nominal gait motions. The control framework is constant between training and deployment to improve sim-to-real transfer.
@@
In contrast, biomechanics research has shown that no single gait is universally optimal across all scenarios within Froude-characterised locomotion [15, 16, 17]. Animals adapt their locomotion by employing nominal gaits such as ambling, trotting, and running [18], while switching to specialised gaits like hopping, pronking, and bounding for off-nominal tasks such as predator evasion or obstacle navigation [19]. Current DRL frameworks fall short of replicating this level of Froude-characterised locomotion versatility. To address this limitation, some approaches have focused on training DRL policies to learn multiple gaits by providing reference motions during training [20, 21, 22, 23], or by learning from policies that specialise in specific gaits [24]. However, these methods remain insufficient when compared to the extensive capabilities observed in animal locomotion, which include:
@@
Although existing DRL frameworks have demonstrated progress in implementing learned gaits, none successfully integrate all three attributes simultaneously. This gap highlights the significant potential of biomechanics-inspired approaches to advance robotic locomotion.
@@
The concept of gait procedural memory, which facilitates the rapid deployment of a range of gaits, is thought to reside within the cerebellum of the animal brain. This region governs the coordination of limb movements for each gait learned by the animal [42, 43]. Similarly, adaptive motion adjustments—crucial for seamless adaptation to off-nominal contact states—are achieved through coordination between the mesencephalic locomotor region (MLR), which oversees locomotion execution [44], and the cerebellum. These adjustments rely on sensory feedback to modify limb movements in response to the animal’s current state [45].
@@ I-A Achieving Adaptive Motion Adjustment with a Diverse Set of Gaits
Figure 3: Locomotion comparison study experiments (a) each policy follows a set of command velocities in $x$ and yaw between 00 to $1$ m/s and $-1$ to $1$ rad/s respectively. During each velocity pair, the commanded gait is cycled through all gaits, switching every $1$ s. This repeated $5$ times over flat, $h_{\text{terr}}^{0}$, to very rough terrain, $h_{\text{terr}}^{3}$ (shown in Figure 2) with the average performance being plotted. A number rather than a magnitude indicates the count of experiments that the policy failed. (b) each policy follows a sinusoidal trajectory in $v_{x}$ and $\omega_{z}$ while switching gaits every $2$ s, which is repeated $5$ times over $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$ terrain. Green highlighted areas indicating transition phases, for which the function and formulation are detailed in Section III-B, and black triangles representing points of failure. (c) policies follow a command of just $v_{x}=0.5$ m/s while using a constant trot gait while encountering rectangular steps with a height of $0.05$ m, where $|\bar{\omega}_{B}^{\text{err}}|$ is the magnitude of the desired orientation error.
@@
As shown in Figure 3(a), on flat terrain, $\pi_{L}^{\text{bio}}$ achieves lower $v_{B}^{\text{err}}$, $c_{\text{avg}}^{\text{err}}$, and $\omega_{B}^{\text{err}}$ than $\pi_{L}^{\text{no}\bm{\beta}{L}}$ and $\pi{L}^{\text{noSE}}$, averaging $15\%$, $21\%$, and $10\%$ lower respectively (excluding failures), with errors increasing at higher velocity commands. This performance gap widens on rough terrain, where $\pi_{L}^{\text{no}\bm{\beta}{L}}$ and $\pi{L}^{\text{noSE}}$ fail frequently—especially at higher speeds and rougher surfaces—while $\pi_{L}^{\text{bio}}$ succeeds in all trials despite being trained only on flat terrain. This highlights its adaptability to unseen environments (see Section III-C for rationale behind omitting rough terrain in $\pi_{L}$ training).
@@
For $\pi_{L}^{\text{no}\bm{\beta}{L}}$, Figure 3(b) shows large spikes in $c_{\text{avg}}^{\text{err}}$ and $\omega_{B}^{\text{err}}$ when encountering $h_{\text{terr}}^{3}$, highlighting its inability to adapt swing foot trajectories and overcome an unrefined solution space. Figure 3(c) further shows sustained instabilities in base height and orientation after contact with steps at $17\%$ of the nominal base height. For $\pi_{L}^{\text{noSE}}$, Figs. 3(b) and 3(c) show that poor reference tracking and stability worsen with velocity command magnitude and time, as it lacks strategies to mitigate error build-up in the SE. With $\pi_{L}^{\text{bio}}$ not experiencing any of these limitations, this explicitly demonstrates the effectiveness of implementing $\bm{\beta}_{L}$ and $\bm{s}$ within $\bm{o}_{L}$; $\pi_{L}^{\text{bio}}$ successfully generalises across gaits and terrain, demonstraing successful instillation of adaptive motion adjustment and gait procedural memory.
Figure 4: Comparative study between each gait and $\pi_{G}^{\text{uni}}$. For terrains $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{2}$, each isolated gait and $\pi_{G}^{\text{uni}}$ are given a velocity command to follow to assess their performance in terms of CoT, $\tau_{\%}$, $W_{\text{ext}}$, and $c_{\text{avg}}^{\text{err}}$; for each individual gait $\pi_{L}^{\text{bio}}$ just run with the gait statically selected and for $\pi_{G}^{\text{uni}}$ it is coupled with $\pi_{L}^{\text{bio}}$ for autonomous optimal gait selection. Additionally at the bottom of these time series plots, the contact state of the feet and the gait that $\pi_{G}^{\text{uni}}$ is also displayed. The average relative performance in terms of these metrics is displayed in the radar plots in the bottom of the figure, where each gait’s performance is normalised to that of the best performer for each metric; the higher the value within the radar plot, the more effectively that metric has been minimised.
@@ I-B Applying Biomechanics Metrics For Optimal Gait Selection
Directly applying biomechanics metrics to instil animal gait transition strategies is unsuitable due to differences between animals and robots and $\pi_{G}$ training requirements. Instead, we use cost of transport (CoT), torque saturation ($\tau_{\%}$), external work ($W_{\text{ext}}$), and foot contact tracking error ($c_{\text{avg}}^{\text{err}}$) to minimise energy use, actuator-structural forces, mechanical work, and instability, respectively. Details and justification of these metrics are in Section III-E.
In accordance with Section III-F, these metrics are unified within the training of the gait selection policy $\pi_{G}^{\text{uni}}$ to instil the strategies animals use for optimal gait selection for exemplary adaptability. To investigate whether $\pi_{G}^{\text{uni}}$ effectively minimises these metrics through gait selection, the results of competing the highly demanding velocity command trajectory presented in Figure 4 and Supplementary Video S3 for $\pi_{G}^{\text{uni}}$ paired with $\pi_{L}^{\text{bio}}$ are collected, along with that for all individual gaits deployable by $\pi_{L}^{\text{bio}}$, for flat terrain and terrain $h_{\text{terr}}^{2}$.
@@ I-C Comparison Between Robot and Animal Gait Selection
Figure 5: Comparison between animal and robot gait selection policy strategies where across all plots both animal and robot locomote at increasing linear forward velocity. (a) The bottom two plots of all robot data indicates the percentage of each gait utilised at that velocity. Magenta, purple and blue shaded regions indicate the transition phases of $\pi_{G}^{\text{uni}}$, $\pi_{G}^{W_{\text{ext}}}$ and animals respectively. This study compares transition strategies of (top left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\text{CoT}}$ to data collected from dogs [35] and horses [46, 33] in terms of CoT, (top right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ to data collected from horses [37] in terms of foot ground reaction forces, (bottom left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{W_{\text{ext}}}$ to horses [33] in terms of external work, and (bottom right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{c^{\text{err}}}$ to opossums and dogs [35]. (b) Mapping the correlation between different all metrics and the average gait ID selected across all velocities for terrains $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$, with the dataset of each terrain consisting of $31$ data points which themselves are an average of $1000$ samples.
When developing metrics to characterise gait transitions in animals, data is collected over intervals of increasing forward velocity on flat terrain [46, 37, 33, 35]. Hence, within Figure 5(a) we took the same approach. We also repeat this experiment with $h_{\text{terr}}^{3}$ to investigate correlations between the metrics and the effects of introducing rough terrain, as presented in Figure 5(b). We also train four additional $\pi_{G}$ policies that individually minimise energy consumption $\pi_{G}^{\text{CoT}}$, actuator-structural forces $\pi_{G}^{\tau_{\%}}$, mechanical work $\pi_{G}^{W_{\text{ext}}}$ and stability $\pi_{G}^{c^{\text{err}}}$, in accordance with Section III-E, to compare their performance with $\pi_{G}^{\text{uni}}$. One unanimous observation across Figure 5(a) is that animals experience a gait transition phase over a range of velocities [35, 46, 47]. This behaviour is reflected in $\pi_{G}^{\text{uni}}$, where we class a transition phase as where no individual gait occupies more than 75% of the gaits used at a specific speed.
@@ I-C1 Energy Expenditure – Cost of Transport
An animal’s gait transition aligns with the CoT-optimal point, $\lambda^{\text{CoT}}$, to minimise energy cost [5, 33], as shown in Figure 5(a). $\pi_{G}^{\text{uni}}$ mirrors this, tracking the lowest-CoT gait with transitions centred on $\lambda^{\text{CoT}}$. In contrast, $\pi_{G}^{\text{CoT}}$ lacks a defined transition phase and switches earlier, due to training on rough terrain where hopping improves CoT (Figure 4). Figure 5(b) further shows $h_{\text{terr}}^{3}$ induces similar CoT distributions but with greater gait variance, as auxiliary gaits become more effective. This highlights that CoT-only training reduces generality and diverges from natural gait selection.

### v3

@@ Abstract
Legged robots must adapt their gait to navigate unpredictable environments, a challenge that animals master with ease. However, most deep reinforcement learning (DRL) approaches to quadruped locomotion rely on a fixed gait, limiting adaptability to changes in terrain and dynamic state. Here we show that integrating three core principles of animal locomotion-gait transition strategies, gait memory and real-time motion adjustments enables a DRL control framework to fluidly switch among multiple gaits and recover from instability, all without external sensing. Our framework is guided by biomechanics-inspired metrics that capture efficiency, stability and system limits, which are unified to inform optimal gait selection. The resulting framework achieves blind zero-shot deployment across diverse, real-world terrains and substantially significantly outperforms baseline controllers. By embedding biological principles into data-driven control, this work marks a step towards robust, efficient and versatile robotic locomotion, highlighting how animal motor intelligence can shape the next generation of adaptive machines.
@@
Originally inspired by the remarkable adaptability of quadruped mammal locomotion, an ability shaped by innate and environmentally induced factors [1, 2, 3], the field of quadruped robotics has invested substantial resources into developing equally proficient locomotion frameworks. At present, the most advanced systems rely on end-to-end deep reinforcement learning (DRL), which involves training a multilayer perceptron [4] capable of navigating diverse environments.
@@
These frameworks demonstrate impressive realisation of terrestrial locomotion skills, which can be classified into two groups (further detailed in Supplementary Section 1); Froude-characterised locomotion (no desired velocity normal to the ground plane and hence upholds the assumptions of the Froude number [5]) such as walking or running in real-world [6, 7], urban [8, 9] and deformable[7] terrains, and Froude-free locomotion (features a desired velocity normal to the ground plane) such as jumping between platforms [10], climbing over obstacles [11], and sure-footedness [12]. Despite these achievements, when it comes to Froude-characterised locomotion, which can account for about $\approx 70-90\%$ daily animal locomotion [13, 14], the adaptability of these frameworks remains constrained, as most systems are limited to deploying a single targeted gait or locomotion strategy.
@@
Figure 1: Instilling the core animal locomotion proficiency attributes within a DRL locomotion framework. From taking an abstracted view of animal locomotion to determine the attributes of proficient locomotion, we instil animal gait transition strategies within the gait selection policy, $\pi_{G}$, gait procedural memory is embedded within the bio-inspired gait scheduler, and adaptive motion adjustments are realised by the locomotion policy, $\pi_{L}$. $\pi_{G}$ has been trained to minimise the animal gait transition metrics applied to the quadruped robot based on the current robot state, $\bm{s}$, and relevant bio-inspired gait scheduler output, $\bm{\beta}_{G}$, to select the optimal gait, $\Gamma^{*}$. The bio-inspired gait scheduler then generates the gait references informed by $\bm{s}$ from encoded high-level gait parameters. The gait references, $\bm{\beta}_{L}$, are then passed to $\pi_{L}$ to inform of any adjustments to the nominal gait motions. The control framework is constant between training and deployment to improve sim-to-real transfer. Regarding acronyms, MLR is, mesencephalic locomotor region, PD is proportional derivative controller, and IMU is inertial measurement unit.
@@
In contrast, biomechanics research has shown that no single gait is universally optimal across all scenarios within Froude-characterised locomotion [15, 16, 17]. Animals adapt their locomotion by employing nominal gaits such as ambling, trotting, and running [18], while switching to specialised gaits such as hopping, pronking and bounding for off-nominal tasks such as predator evasion or obstacle navigation [19]. Current DRL frameworks fall short of replicating this level of Froude-characterised locomotion versatility. To address this limitation, some approaches have focused on training DRL policies to learn multiple gaits by providing reference motions during training [20, 21, 22, 23], or by learning from policies that specialise in specific gaits [24]. However, these methods remain insufficient when compared to the extensive capabilities observed in animal locomotion, which include:
@@
Although existing DRL frameworks have demonstrated progress in implementing learned gaits, none successfully integrate all three attributes simultaneously. This gap highlights the considerable potential of biomechanics-inspired approaches to advance robotic locomotion.
@@
The concept of gait procedural memory, which facilitates the rapid deployment of a range of gaits, is thought to reside within the cerebellum of the animal brain. This region governs the coordination of limb movements for each gait learned by the animal [42, 43]. Similarly, adaptive motion adjustments, crucial for seamless adaptation to off-nominal contact states, are achieved through coordination between the mesencephalic locomotor region (MLR), which oversees locomotion execution [44], and the cerebellum. These adjustments rely on sensory feedback to modify limb movements in response to the animal’s current state [45].
@@ I-A Achieving Adaptive Motion Adjustment with a Diverse Set of Gaits
Figure 3: Locomotion comparison study experiments (a) each policy follows a set of command velocities in $x$ and yaw between 00 to $1$ m/s and $-1$ to $1$ rad/s respectively. During each velocity pair, the commanded gait is cycled through all gaits, switching every $1$ s. This repeated $5$ times over flat, $h_{\text{terr}}^{0}$, to very rough terrain, $h_{\text{terr}}^{3}$ (shown in Figure 2) with the average performance being plotted. A number rather than a magnitude indicates the count of experiments that the policy failed. (b) each policy follows a sinusoidal trajectory in $x$, $v_{x}$, and in yaw, $\omega_{z}$, while switching gaits every $2$ s, which is repeated $5$ times over $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$ terrain. Green highlighted areas indicating transition phases, for which the function and formulation are detailed in Section III-B, and black triangles representing points of failure. (c) policies follow a command of just $v_{x}=0.5$ m/s while using a constant trot gait while encountering rectangular steps with a height of $0.05$ m, where $|\bar{\omega}_{B}^{\text{err}}|$ is the magnitude of the desired orientation error. FL is front left, FR is front right, RL is rear left, and RR is rear right in reference to the robot’s feet.
@@
As shown in Figure 3a, on flat terrain, $\pi_{L}^{\text{bio}}$ achieves lower $v_{B}^{\text{err}}$, $c_{\text{avg}}^{\text{err}}$, and $\omega_{B}^{\text{err}}$ than $\pi_{L}^{\text{no}\bm{\beta}{L}}$ and $\pi_{L}^{\text{noSE}}$, averaging $15\%$, $21\%$, and $10\%$ lower respectively (excluding failures), with errors increasing at higher velocity commands. This performance gap widens on rough terrain, where $\pi_{L}^{\text{no}\bm{\beta}{L}}$ and $\pi_{L}^{\text{noSE}}$ fail frequently, especially at higher speeds and rougher surfaces, while $\pi_{L}^{\text{bio}}$ succeeds in all trials despite being trained only on flat terrain. This highlights its adaptability to unseen environments (see Section III-C for rationale behind omitting rough terrain in $\pi_{L}$ training).
@@
For $\pi_{L}^{\text{no}\bm{\beta}{L}}$, Figure 3b shows large spikes in $c_{\text{avg}}^{\text{err}}$ and $\omega_{B}^{\text{err}}$ when encountering $h_{\text{terr}}^{3}$, highlighting its inability to adapt swing foot trajectories and overcome an unrefined solution space. Figure 3c further shows sustained instabilities in base height and orientation after contact with steps at $17\%$ of the nominal base height. For $\pi_{L}^{\text{noSE}}$, Figs. 3b and 3c show that poor reference tracking and stability worsen with velocity command magnitude and time, as it lacks strategies to mitigate error build-up in the SE. Since $\pi_{L}^{\text{bio}}$ does not experience any of these limitations, this explicitly demonstrates the effectiveness of implementing $\bm{\beta}_{L}$ and $\bm{s}$ within $\bm{o}_{L}$; $\pi_{L}^{\text{bio}}$ successfully generalises across gaits and terrain, demonstraing successful instillation of adaptive motion adjustment and gait procedural memory.
Figure 4: Comparative study between each gait and $\pi_{G}^{\text{uni}}$. For terrains $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{2}$, each isolated gait and $\pi_{G}^{\text{uni}}$ are given a velocity command to follow to assess their performance in terms of CoT, $\tau_{\%}$, $W_{\text{ext}}$, and $c_{\text{avg}}^{\text{err}}$; for each individual gait $\pi_{L}^{\text{bio}}$ just run with the gait statically selected and for $\pi_{G}^{\text{uni}}$ it is coupled with $\pi_{L}^{\text{bio}}$ for autonomous optimal gait selection. Additionally at the bottom of these time series plots, the contact state of the feet and the gait that $\pi_{G}^{\text{uni}}$ realises is also shown. The average relative performance in terms of these metrics is displayed in the radar plots at the bottom of the figure, where each gait’s performance is normalised to that of the best performer for each metric; the higher the value within the radar plot, the more effectively that metric has been minimised.
@@ I-B Applying Biomechanics Metrics For Optimal Gait Selection
Directly applying biomechanics metrics to instil animal gait transition strategies is unsuitable due to differences between animals and robots, as well as $\pi_{G}$ training requirements. Instead, we use cost of transport (CoT), torque saturation ($\tau_{\%}$), external work ($W_{\text{ext}}$), and foot contact tracking error ($c_{\text{avg}}^{\text{err}}$) to minimise energy use, actuator-structural forces, mechanical work, and instability, respectively. Details and justification of these metrics are in Section III-E.
In accordance with Section III-F, these metrics are unified within the training of the gait selection policy $\pi_{G}^{\text{uni}}$ to instil the strategies animals use for optimal gait selection to acheive exemplary adaptability. To investigate whether $\pi_{G}^{\text{uni}}$ effectively minimises these metrics through gait selection, the results of competing the highly demanding velocity command trajectory presented in Figure 4 and Supplementary Video S3 for $\pi_{G}^{\text{uni}}$ paired with $\pi_{L}^{\text{bio}}$ are collected, along with that for all individual gaits deployable by $\pi_{L}^{\text{bio}}$, for flat terrain and terrain $h_{\text{terr}}^{2}$.
@@ I-C Comparison Between Robot and Animal Gait Selection
Figure 5: Comparison between animal and robot gait selection policy strategies during increasing linear velocity. (a) The bottom two plots of all robot data indicate the percentage of each gait utilised at that velocity. Magenta, purple and blue shaded regions indicate the transition phases of $\pi_{G}^{\text{uni}}$, $\pi_{G}^{W_{\text{ext}}}$ and animals respectively. This study compares transition strategies of (top left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\text{CoT}}$ to data collected from dogs [35] and horses [46, 33] in terms of CoT, (top right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ to data collected from horses [37] in terms of foot ground reaction forces, (bottom left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{W_{\text{ext}}}$ to horses [33] in terms of external work, and (bottom right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{c^{\text{err}}}$ to opossums and dogs [35]. (b) Mapping the correlation between metrics and the average gait ID selected across velocities for terrains $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$, with the dataset of each terrain consisting of $31$ data points averaged over $1000$ samples.
When developing metrics to characterise gait transitions in animals, data is collected over intervals of increasing forward velocity on flat terrain [46, 37, 33, 35]. Hence, within Figure 5a we took the same approach. We also repeat this experiment with $h_{\text{terr}}^{3}$ to investigate correlations between the metrics and the effects of introducing rough terrain, as presented in Figure 5b. Additionally, we train four further $\pi_{G}$ policies that individually minimise energy consumption $\pi_{G}^{\text{CoT}}$, actuator-structural forces $\pi_{G}^{\tau_{\%}}$, mechanical work $\pi_{G}^{W_{\text{ext}}}$ and stability $\pi_{G}^{c^{\text{err}}}$, in accordance with Section III-E, to compare their performance with $\pi_{G}^{\text{uni}}$. One unanimous observation across Figure 5a is that animals experience a gait transition phase over a range of velocities [35, 46, 47]. This behaviour is reflected in $\pi_{G}^{\text{uni}}$, where we class a transition phase as where no individual gait occupies more than 75% of the gaits used at a specific speed.
@@ I-C1 Energy Expenditure – Cost of Transport
An animal’s gait transition aligns with the CoT-optimal point, $\lambda^{\text{CoT}}$, to minimise energy cost [5, 33], as shown in Figure 5a. $\pi_{G}^{\text{uni}}$ mirrors this, tracking the lowest-CoT gait with transitions centred on $\lambda^{\text{CoT}}$. In contrast, $\pi_{G}^{\text{CoT}}$ lacks a defined transition phase and switches earlier, due to training on rough terrain where hopping improves CoT (Figure 4). Figure 5b further shows $h_{\text{terr}}^{3}$ induces similar CoT distributions but with greater gait variance, as auxiliary gaits become more effective. This highlights that CoT-only training reduces generality and diverges from natural gait selection.